OpenStudy (landon34):

Three-fourths was subtracted from a number, and then 5 1/4 was added to the result. the result was -7. what was the original number?

OpenStudy (landon34):

@jabez177

OpenStudy (landon34):

ok can you help me out?

OpenStudy (just_one_last_goodbye):

I really dont know sorry ;-;

OpenStudy (landon34):

oh.

OpenStudy (landon34):

could you help me out?

Elsa213 (elsa213):

There are no answer choices?

OpenStudy (landon34):

oh yeah, hold on.

OpenStudy (landon34):

-3 1/2 3 1/2 -11 1/2 2 1/2

Elsa213 (elsa213):

What do you think the correct answer is?

OpenStudy (landon34):

is it -11 1/2?

OpenStudy (landon34):

srry Im not good with fractions

Elsa213 (elsa213):

hrmm..... @pooja195

pooja195 (pooja195):

\[\huge~\rm~\frac{ 3 }{ 4 }-x+5 \frac{ 1 }{ 4 }=-7\] Solve for x

OpenStudy (landon34):

lemme write that down.

OpenStudy (landon34):

ok.

OpenStudy (landon34):

yes, I was right, it is 11 1/2

OpenStudy (landon34):

thank you for the help!

OpenStudy (landon34):

why is everybody still on here?

Elsa213 (elsa213):

I dun know

OpenStudy (landon34):

oh yeah, forgot to close.

OpenStudy (anonymous):

@Thadds2003

OpenStudy (landon34):

thanks for reminding me.

OpenStudy (anonymous):

because you bumped it up

OpenStudy (landon34):

oops.